OpenStudy (anonymous):

What number must you add to complete the square? x2 + 12x = -9

OpenStudy (anonymous):

is that supposed to be 2x or x^2?

OpenStudy (gorv):

add 9 both side

OpenStudy (gorv):

x2+12x+9=0

OpenStudy (gorv):

now seprate

OpenStudy (anonymous):

use mathway .com all you do is enter the equation and you can choose how you want it to solve for you and it will give you the answer. really useful and simple and its in the app store retricewell, not sure about google play

OpenStudy (gorv):

x2+2*6*x+9+27-27=0 x^2+2*6*x+36-27

OpenStudy (gorv):

(x+6)^2-27=0

OpenStudy (gorv):

add 27 both side

OpenStudy (gorv):

(x+6)^2=27

OpenStudy (gorv):

27 we need to add